# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v2

@@ 7.1 Strategy of reduction
Therefore, the transformed model is equivalent to the standard sequence model with uncorrelated noise in Equation˜4 and the estimation is equivalent to the estimation therein. Consequently, the ESD and span profile for the model in Equation˜14 can be naturally defined using the original definitions for the transformed model. Specifically, for the correlated-noise model, we define the ESD w.r.t. $\Sigma_{\xi}$ as
@@
This strategy of reducing a complex model to a simple model will be used in our analysis of the linear model and in RKHS regression in Section˜8.
@@ 7.2 Linear model
where $\mathbf{Y}\in\mathbb{R}^{n}$ is the vector of observations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the fixed-design matrix of rank $r\leq\min(n,p)$, $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the unknown vector of true coefficients, and $\bm{\epsilon}\in\mathbb{R}^{n}$ is the noise vector. We assume the components of $\bm{\epsilon}$ are uncorrelated with mean zero and variance $\sigma_{0}^{2}$. For this model, we consider the loss (in-sample prediction error) $\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\frac{1}{n}\|\mathbf{X}\left(\widehat{\bm{\beta}}-\bm{\beta}^{*}\right)\|^{2}$ and risk $\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})$. For random design linear regression, we treat it as a special case of RKHS regression and discuss it in Section˜8.
@@
Multiplying the model in Equation˜15 by $\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}$, we obtain a $r$-dimensional transformed model:
@@
The transformed model in Equation˜17 is analogous to the sequence model in Equation˜4, where the signal is $\bm{\theta}^{*}$ and the noise variance for each component is $\sigma^{2}$. The “spectrum” relevant to this problem is derived from the singular values of $\mathbf{X}$. Specifically, we define the eigenvalues as $\lambda_{j}=s_{j}^{2}$ for $j=1,\dots,r$, and $\lambda_{j}=0$ for $j>r$. Let $\{\pi_{k}\}_{k=1}^{r}$ denote the indices corresponding to the eigenvalues sorted in descending order, $\lambda_{\pi_{1}}\geq\lambda_{\pi_{2}}\geq\dots\geq\lambda_{\pi_{r}}>0$.
For any estimator $\widehat{\bm{\beta}}$ for the linear model Equation˜15, define $\widehat{\bm{\theta}}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}$. We can then write the prediction risk as $n^{-1}\mathbb{E}\|\mathbf{X}\widehat{\bm{\beta}}-\mathbf{X}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\widehat{\bm{\beta}}-\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\widehat{\bm{\theta}}-\bm{\theta}^{*}\|^{2}$. Conversely, given an estimator $\widehat{\bm{\theta}}$ for the sequence model Equation˜17, we can define $\widetilde{\bm{\beta}}=\mathbf{V}\mathbf{S}^{\dagger}\widehat{\bm{\theta}}$ where $\mathbf{S}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{1/s_{j}\}_{j\in[r]}$. It is easy to check that $\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widetilde{\bm{\beta}}=\widehat{\bm{\theta}}$. Therefore, we establish an equivalence between the model Equation˜17 and the model Equation˜15.
The usual ridge regression estimator for the linear model Equation˜15 is given by
@@ 7.3 ESD for linear models
Suppose the SVD of the design matrix $\mathbf{X}$ is given in Equation˜16. The Effective Span Dimension (ESD) of $\bm{\beta}^{*}$ with respect to the design $\mathbf{X}$ and the per component variance $\sigma_{0}^{2}/n$ is defined as
@@
The Principal Component Regression (PCR) estimator for $\bm{\beta}^{*}$ corresponds to the Principal Component (PC) estimator in the transformed space Equation˜17. Specifically, for any $k\in[r]$, define $\widehat{\bm{\beta}}^{\operatorname{PC},k}=\frac{1}{\sqrt{n}}\mathbf{V}\mathbf{S}_{k}^{\dagger}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, where $\mathbf{S}_{k}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{\frac{1}{s_{j}}\mathbf{1}_{\{s_{j}\geq s_{\pi_{k}}\}}\}$. In the $\mathbf{Z}$ space, this means
@@
where $d^{\dagger}=d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})$ is the ESD defined in Definition˜7.1.
Proposition˜7.2 directly follows from Theorem˜3.2 and its proof is omitted. This result shows that the optimal prediction risk for PCR is determined by the ESD $d^{\dagger}$, which measures the effective number of principal components needed to balance the bias-variance trade-off.
@@
Suppose $K\leq r$. For the linear model Equation˜15 with noise variance $\sigma_{0}^{2}$, the minimax prediction risk over the class $\mathcal{B}_{K}^{(n)}$ defined in Equation˜19 satisfies:
@@
The proof of Theorem˜7.3 is essentially the same as that of Theorem˜3.3 and is omitted.
@@ 7.4 Numerical illustration
Figure 3 plots $d^{\dagger}(\alpha)$ (solid) and the rescaled oracle risk defined as $n\mathcal{R}_{*}(\alpha)/\sigma_{0}^{2}$ (dashed) against $\alpha$. The two curves coincide over the entire path, which empirically verifies the bound in Proposition˜7.2 that $\mathcal{R}_{*}(\alpha)\asymp\frac{\sigma_{0}^{2}}{n}\,d^{\dagger}(\alpha)$. As $\alpha$ grows, the diagonal stretch $\mathbf{A}(\alpha)$ shifts signal energy towards directions that carry smaller singular values. This raises $d^{\dagger}$, and consequently, the achievable risk.
@@ 8 Extension to RKHS Regression
This section extends the concepts of Effective Span Dimension (ESD) and span profile, developed in Section˜3, to the setting of RKHS regression. Since our goal is to develop a population level complexity measure, we focus on the simple case where the eigenfunctions of the kernel are fully known and computable, and leave a thorough analysis in future studies.
@@ 8.1 RKHS regression
The expression in Equation˜21 suggests that we can equivalently estimate each $\theta_{j}^{*}$ separately using the transformed observation $z_{j}=n^{-1}\sum_{i}y_{i}\psi_{j}(x_{i})$ for any $j\geq 1$ as introduced in Equation˜2.
@@ 8.2 ESD for RKHS Regression
˜8.1 is very mild: for compact $\mathcal{X}$, if $f$ is continuous, then $f$ is bounded.
@@
where $\bm{\theta}^{*}=\{\theta_{j}^{*}\}_{j\geq 1}$ and $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)$ is defined as in Equation˜10.
@@
Let $\hat{f}_{k}^{\operatorname{PC}}$ be the KPCPE estimator defined in Equation˜22. Let $\mathcal{R}_{*}^{\operatorname{PC}}=\min_{k\geq 1}\mathcal{R}(\hat{f}_{k}^{\operatorname{PC}};f^{*})$. Under Assumption 8.1, it holds that:
@@
We comment that Zhang et al. [50] have established the minimax optimality of the well-tuned PC estimator. Proposition˜8.3 suggests that the risk of the well-tuned PC estimator scales as $d^{\dagger}/n$. Therefore, we essentially express the minimax rate therein using the ESD without reliance on the classical eigen-decay conditions or source conditions.
@@
Combining Theorem˜8.5 and Proposition˜8.3, the optimally tuned KPCPE estimator is minimax rate optimal over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$, with rate $\sigma_{0}^{2}K/n$.
@@
where $\bar{\sigma}^{2}=\sigma_{0}^{2}(1+C_{0}^{2})$. For a sample $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from the model in Equation˜20 and any estimator $\hat{f}$, we aim to determine the optimal convergence rate of the following minimax risk:
@@
Suppose ˜4.1 holds for a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. Furthermore, suppose $\mathbf{k}$ is $(K_{n},n)$-regular for all $n\geq n_{0}$. If $\{(x_{i},y_{i})\}_{i=1}^{n}$ is drawn from the model in Equation˜20, it holds that
@@
˜8.4 is a mild condition. The following is an example where we use Theorem˜8.6 to recover the minimax convergence rate derived under the classical polynomial eigen-decay conditions and source conditions.
@@
regardless the regularization parameter. In that case, we may modify the definition of ESD by adding the systematic bias into the summation to the sum of squared tails.
Specifically, for a PSD kernel $\mathbf{k}$ of rank $d<\infty$, we modify the definition of ESD in Definition˜8.2 as
@@
Again, $\sigma^{2}\bar{d}^{\dagger}$ characterizes the risk of the well-tuned PC estimator as in Proposition˜8.3.
@@ 8.3 Connection to Random Design Linear Regression
Therefore, with respect to the kernel $\mathbf{k}$ and the basis $\{\psi_{j}\}$, we define the ESD exactly as in Definition˜8.2 using the coefficients $\{\theta^{*}_{j}\}$ and eigenvalues $\{\lambda_{j}\}$.
@@ 8.4 Numerical Illustration
This section provides numerical validation of the relationship between the ESD and the optimally tuned KPCPE risk, mirroring the setup for linear models in Section˜7.4. We use the cosine basis eigenfunctions $\psi_{j}(x)=\sqrt{2}\cos(2\pi jx)$ on the domain $[0,1]$ with inputs sampled as $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Unif}[0,1]$. The sample size is fixed at $n=400$, and for numerical purpose, we consider the first $J=800$ eigenfunctions. The noise variance is set as $\sigma_{0}^{2}=1$.
@@ 9.2 Pathwise ESD for Learned Kernels
Section˜5 analyzes eigenvalue learning because OP-GF admits tractable dynamics under a fixed eigenbasis. This is a limitation of that specific analysis, not of the ESD concept. In fact, ESD applies to general representation learning. We illustrate how decreases in ESD explain minimax risk reduction for learned kernels, whether adaptation acts through eigenvalues, eigenfunctions, or both.
@@
where we have followed Definition˜8.2 to define $d^{\dagger}(\sigma^{2};f^{*},\mathbf{k}_{t})$ as the ESD of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$ using $\theta_{j}^{*,(t)}=\langle f^{*},\psi_{j}^{(t)}\rangle$ and $\sigma^{2}=n^{-1}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$.
@@
The estimated function at time $t$ is given by $f_{t}(x)=w(t)^{\top}\mathbf{A}(t)x$, where $\mathbf{A}(t):=\mathbf{W}_{D}(t)\cdots\mathbf{W}_{1}(t)$. We form the learned kernel $\mathbf{k}_{t}(x,x^{\prime})=\langle\mathbf{A}(t)x,\mathbf{A}(t)x^{\prime}\rangle=x^{\top}\mathbf{G}_{t}x^{\prime}$, where $G_{t}:=\mathbf{A}(t)^{\top}\mathbf{A}(t)$. We then follow the derivation in Section˜8.3 and define the ESD $d^{\dagger}(t)$ of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$. Since $\|X\|_{\infty}=1$ $\mu$-a.s., we have $\|f^{*}\|_{\infty}^{2}=\|\beta^{*}\|_{1}^{2}$; this is used in computing the effective noise level.

### v3

@@ 8.1 Strategy of reduction
Therefore, the transformed model is equivalent to the standard sequence model with uncorrelated noise in Equation˜4 and the estimation is equivalent to the estimation therein. Consequently, the ESD and span profile for the model in Equation˜15 can be naturally defined using the original definitions for the transformed model. Specifically, for the correlated-noise model, we define the ESD w.r.t. $\Sigma_{\xi}$ as
@@
This strategy of reducing a complex model to a simple model will be used in our analysis of the linear model and in RKHS regression in Section˜9.
@@ 8.2 Linear model
where $\mathbf{Y}\in\mathbb{R}^{n}$ is the vector of observations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the fixed-design matrix of rank $r\leq\min(n,p)$, $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the unknown vector of true coefficients, and $\bm{\epsilon}\in\mathbb{R}^{n}$ is the noise vector. We assume the components of $\bm{\epsilon}$ are uncorrelated with mean zero and variance $\sigma_{0}^{2}$. For this model, we consider the loss (in-sample prediction error) $\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\frac{1}{n}\|\mathbf{X}\left(\widehat{\bm{\beta}}-\bm{\beta}^{*}\right)\|^{2}$ and risk $\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})$. For random design linear regression, we treat it as a special case of RKHS regression and discuss it in Section˜9.
@@
Multiplying the model in Equation˜16 by $\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}$, we obtain a $r$-dimensional transformed model:
@@
The transformed model in Equation˜18 is analogous to the sequence model in Equation˜4, where the signal is $\bm{\theta}^{*}$ and the noise variance for each component is $\sigma^{2}$. The “spectrum” relevant to this problem is derived from the singular values of $\mathbf{X}$. Specifically, we define the eigenvalues as $\lambda_{j}=s_{j}^{2}$ for $j=1,\dots,r$, and $\lambda_{j}=0$ for $j>r$. Let $\{\pi_{k}\}_{k=1}^{r}$ denote the indices corresponding to the eigenvalues sorted in descending order, $\lambda_{\pi_{1}}\geq\lambda_{\pi_{2}}\geq\dots\geq\lambda_{\pi_{r}}>0$.
For any estimator $\widehat{\bm{\beta}}$ for the linear model Equation˜16, define $\widehat{\bm{\theta}}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}$. We can then write the prediction risk as $n^{-1}\mathbb{E}\|\mathbf{X}\widehat{\bm{\beta}}-\mathbf{X}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\widehat{\bm{\beta}}-\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\widehat{\bm{\theta}}-\bm{\theta}^{*}\|^{2}$. Conversely, given an estimator $\widehat{\bm{\theta}}$ for the sequence model Equation˜18, we can define $\widetilde{\bm{\beta}}=\mathbf{V}\mathbf{S}^{\dagger}\widehat{\bm{\theta}}$ where $\mathbf{S}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{1/s_{j}\}_{j\in[r]}$. It is easy to check that $\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widetilde{\bm{\beta}}=\widehat{\bm{\theta}}$. Therefore, we establish an equivalence between the model Equation˜18 and the model Equation˜16.
The usual ridge regression estimator for the linear model Equation˜16 is given by
@@ 8.3 ESD for linear models
Suppose the SVD of the design matrix $\mathbf{X}$ is given in Equation˜17. The Effective Span Dimension (ESD) of $\bm{\beta}^{*}$ with respect to the design $\mathbf{X}$ and the per component variance $\sigma_{0}^{2}/n$ is defined as
@@
The Principal Component Regression (PCR) estimator for $\bm{\beta}^{*}$ corresponds to the Principal Component (PC) estimator in the transformed space Equation˜18. Specifically, for any $k\in[r]$, define $\widehat{\bm{\beta}}^{\operatorname{PC},k}=\frac{1}{\sqrt{n}}\mathbf{V}\mathbf{S}_{k}^{\dagger}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, where $\mathbf{S}_{k}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{\frac{1}{s_{j}}\mathbf{1}_{\{s_{j}\geq s_{\pi_{k}}\}}\}$. In the $\mathbf{Z}$ space, this means
@@
where $d^{\dagger}=d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})$ is the ESD defined in Definition˜8.1.
Proposition˜8.2 directly follows from Theorem˜3.2 and its proof is omitted. This result shows that the optimal prediction risk for PCR is determined by the ESD $d^{\dagger}$, which measures the effective number of principal components needed to balance the bias-variance trade-off.
@@
Suppose $K\leq r$. For the linear model Equation˜16 with noise variance $\sigma_{0}^{2}$, the minimax prediction risk over the class $\mathcal{B}_{K}^{(n)}$ defined in Equation˜20 satisfies:
@@
The proof of Theorem˜8.3 is essentially the same as that of Theorem˜3.3 and is omitted.
@@ 8.4 Numerical illustration
Figure 3 plots $d^{\dagger}(\alpha)$ (solid) and the rescaled oracle risk defined as $n\mathcal{R}_{*}(\alpha)/\sigma_{0}^{2}$ (dashed) against $\alpha$. The two curves coincide over the entire path, which empirically verifies the bound in Proposition˜8.2 that $\mathcal{R}_{*}(\alpha)\asymp\frac{\sigma_{0}^{2}}{n}\,d^{\dagger}(\alpha)$. As $\alpha$ grows, the diagonal stretch $\mathbf{A}(\alpha)$ shifts signal energy towards directions that carry smaller singular values. This raises $d^{\dagger}$, and consequently, the achievable risk.
@@ 9 Extension to RKHS Regression
This section extends the concepts of Effective Span Dimension (ESD) and span profile, developed in Section˜3, to the setting of RKHS regression. Since our goal is to develop a population-level complexity measure, we focus on the simple case where the eigenfunctions of the kernel are fully known and computable, and leave a thorough analysis for future studies.
@@ 9.1 RKHS regression
The expression in Equation˜22 suggests that we can equivalently estimate each $\theta_{j}^{*}$ separately using the transformed observation $z_{j}=n^{-1}\sum_{i}y_{i}\psi_{j}(x_{i})$ for any $j\geq 1$ as introduced in Equation˜2.
@@ 9.2 ESD for RKHS Regression
˜9.1 is very mild: for compact $\mathcal{X}$, if $f$ is continuous, then $f$ is bounded.
@@
where $\bm{\theta}^{*}=\{\theta_{j}^{*}\}_{j\geq 1}$ and $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)$ is defined as in Equation˜9.
@@
Let $\hat{f}_{k}^{\operatorname{PC}}$ be the KPCPE estimator defined in Equation˜23. Let $\mathcal{R}_{*}^{\operatorname{PC}}=\min_{k\geq 1}\mathcal{R}(\hat{f}_{k}^{\operatorname{PC}};f^{*})$. Under Assumption 9.1, it holds that:
@@
We comment that Zhang et al. [51] have established the minimax optimality of the well-tuned PC estimator. Proposition˜9.3 suggests that the risk of the well-tuned PC estimator scales as $d^{\dagger}/n$. Therefore, we essentially express the minimax rate therein using the ESD without reliance on the classical eigen-decay conditions or source conditions.
@@
Combining Theorem˜9.5 and Proposition˜9.3, the optimally tuned KPCPE estimator is minimax rate optimal over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$, with rate $\sigma_{0}^{2}K/n$.
@@
where $\bar{\sigma}^{2}=\sigma_{0}^{2}(1+C_{0}^{2})$. For a sample $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from the model in Equation˜21 and any estimator $\hat{f}$, we aim to determine the optimal convergence rate of the following minimax risk:
@@
Suppose ˜4.1 holds for a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. Furthermore, suppose $\mathbf{k}$ is $(K_{n},n)$-regular for all $n\geq n_{0}$. If $\{(x_{i},y_{i})\}_{i=1}^{n}$ is drawn from the model in Equation˜21, it holds that
@@
˜9.4 is a mild condition. The following is an example where we use Theorem˜9.6 to recover the minimax convergence rate derived under the classical polynomial eigen-decay conditions and source conditions.
@@
regardless of the regularization parameter. In that case, we may modify the definition of ESD by adding the systematic bias into the summation to the sum of squared tails.
Specifically, for a PSD kernel $\mathbf{k}$ of rank $d<\infty$, we modify the definition of ESD in Definition˜9.2 as
@@
Again, $\sigma^{2}\bar{d}^{\dagger}$ characterizes the risk of the well-tuned PC estimator as in Proposition˜9.3.
@@ 9.3 Connection to Random Design Linear Regression
Therefore, with respect to the kernel $\mathbf{k}$ and the basis $\{\psi_{j}\}$, we define the ESD exactly as in Definition˜9.2 using the coefficients $\{\theta^{*}_{j}\}$ and eigenvalues $\{\lambda_{j}\}$.
@@ 9.4 Numerical Illustration
This section provides numerical validation of the relationship between the ESD and the optimally tuned KPCPE risk, mirroring the setup for linear models in Section˜8.4. We use the cosine basis eigenfunctions $\psi_{j}(x)=\sqrt{2}\cos(2\pi jx)$ on the domain $[0,1]$ with inputs sampled as $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Unif}[0,1]$. The sample size is fixed at $n=400$, and for numerical purpose, we consider the first $J=800$ eigenfunctions. The noise variance is set as $\sigma_{0}^{2}=1$.
@@ 10.2 Pathwise ESD for Learned Kernels
Section˜6 analyzes eigenvalue learning because OP-GF admits tractable dynamics under a fixed eigenbasis. This is a limitation of that specific analysis, not of the ESD concept. In fact, ESD applies to general representation learning. We illustrate how decreases in ESD explain minimax risk reduction for learned kernels, whether adaptation acts through eigenvalues, eigenfunctions, or both.
@@
where we have followed Definition˜9.2 to define $d^{\dagger}(\sigma^{2};f^{*},\mathbf{k}_{t})$ as the ESD of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$ using $\theta_{j}^{*,(t)}=\langle f^{*},\psi_{j}^{(t)}\rangle$ and $\sigma^{2}=n^{-1}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$.
@@
The estimated function at time $t$ is given by $f_{t}(x)=w(t)^{\top}\mathbf{A}(t)x$, where $\mathbf{A}(t):=\mathbf{W}_{D}(t)\cdots\mathbf{W}_{1}(t)$. We form the learned kernel $\mathbf{k}_{t}(x,x^{\prime})=\langle\mathbf{A}(t)x,\mathbf{A}(t)x^{\prime}\rangle=x^{\top}\mathbf{G}_{t}x^{\prime}$, where $G_{t}:=\mathbf{A}(t)^{\top}\mathbf{A}(t)$. We then follow the derivation in Section˜9.3 and define the ESD $d^{\dagger}(t)$ of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$. Since $\|X\|_{\infty}=1$ $\mu$-a.s., we have $\|f^{*}\|_{\infty}^{2}=\|\beta^{*}\|_{1}^{2}$; this is used in computing the effective noise level.